# FMPlug: Plug-In Foundation Flow-Matching Priors for Inverse Problems

### v1

@@ 3.2 Gaussianity regularization
If $\bm{z}_{0}\sim\mathcal{N}(\bm{0},\bm{I})$, $\norm{\bm{z}_{0}}_{2}^{2}\sim\chi^{2}(d)$ and the negative log-likelihood is $h(\bm{z}_{0})=-(d/2-1)\log\norm{\bm{z}_{0}}_{2}^{2}+\norm{\bm{z}_{0}}_{2}^{2}/2+C$ for some constant $C$ independent of $\bm{z}_{0}$. [19] promotes the Gaussianity of $\bm{z}_{0}$ by regularizing $h(\bm{z}_{0})$. While $h(\bm{z}_{0})$ is minimized at any $\bm{z}_{0}$ satisfies $\norm{\bm{z}_{0}}_{2}^{2}=d-2$, away from this value the function changes slowly; see Fig. 1. This is problematic, as $\norm{\bm{z}_{0}}_{2}^{2}$ should concentrate sharply around $d$ due to concentration of measure and thus only functions that blow up quickly away from the $\norm{\bm{z}_{0}}_{2}^{2}=d$ level can effectively promote the Gaussianity of $\bm{z}_{0}$.
@@
We are not the first to use the more precise spherical constraint or regularization in promoting Gaussianity in the DM/FM literature; see, e.g., [24].
@@ 4 Experiment
DIV2K [25] AFHQ [26]
@@
DIV2K [25] AFHQ [26]
@@
Datasets, tasks, and evaluation metrics  Following [17], we use two datasets: 252525 DIV2K  [25] validation images and 252525 AFHQ cat  [26] images. We set the image resolution to $\times 512$ 512 by resizing and cropping the original. We consider two tasks: i) 4$\times$ super-resolution (SR) from $\times 128$ 128 to $\times 512$ 512; ii) Gaussian deblurring (GD) with a kernel size of 616161 and standard deviation of 3.03.03.0. Both tasks are added with Gaussian noise $\sigma=0.03$. For metrics, we use PSNR and SSIM for pixel-level difference, and LPIPS for perceptual difference, and CLIPIQA, a no-reference quality metric.
@@
Fig. 2: Qualitative results for 4×4$\times$4 × super-resolution on the AFHQ dataset.
@@
Fig. 3: Qualitative results for 4×4$\times$4 × super-resolution on the DIV2K dataset.

### v2

@@ 3.2 Gaussianity regularization
If $\bm{z}_{0}\sim\mathcal{N}(\bm{0},\bm{I})$, $\norm{\bm{z}_{0}}_{2}^{2}\sim\chi^{2}(d)$ and the negative log-likelihood is $h(\bm{z}_{0})=-(d/2-1)\log\norm{\bm{z}_{0}}_{2}^{2}+\norm{\bm{z}_{0}}_{2}^{2}/2+C$ for some constant $C$ independent of $\bm{z}_{0}$. [ben2024d] promotes the Gaussianity of $\bm{z}_{0}$ by regularizing $h(\bm{z}_{0})$. While $h(\bm{z}_{0})$ is minimized at any $\bm{z}_{0}$ satisfies $\norm{\bm{z}_{0}}_{2}^{2}=d-2$, away from this value the function changes slowly; see Fig. 1. This is problematic, as $\norm{\bm{z}_{0}}_{2}^{2}$ should concentrate sharply around $d$ due to concentration of measure and thus only functions that blow up quickly away from the $\norm{\bm{z}_{0}}_{2}^{2}=d$ level can effectively promote the Gaussianity of $\bm{z}_{0}$.
@@
We are not the first to use the more precise spherical constraint or regularization in promoting Gaussianity in the DM/FM literature; see, e.g., [yang2024guidancesphericalgaussianconstraint].
@@ 4 Experiment
DIV2K [Agustsson_2017_CVPR_Workshops] AFHQ [choi2020starganv2]
@@
DIV2K [Agustsson_2017_CVPR_Workshops] AFHQ [choi2020starganv2]
@@
Competing methods  We compare our FMPlug (W: warm-up, R: regularized, Number of Function Evaluations (NFE) =3=3) with Deep Image Prior [Ulyanov_2020] (an untrained image prior), D-Flow ($\text{NFE}=3$) [ben2024d] (a SOTA plug-in method), FlowDPS ($\text{NFE}=28$) [kim2025flowdpsflowdrivenposteriorsampling] (a SOTA interleaving method) and FlowChef ($\text{NFE}=28$) [patel2024steering] (another SOTA interleaving method). For a fair comparison, we use the pretrained SD3 model [esser2024scalingrectifiedflowtransformers] as the backbone for all methods that require pretrained priors, i.e., foundation priors.
@@
Fig. 2: Qualitative results for $4\times$ super-resolution on the AFHQ dataset.
@@
Fig. 3: Qualitative results for $4\times$ super-resolution on the DIV2K dataset.